1. Compare and contrast the three methods of measuring the national income statistics.
  2. With examples discuss the reasons for unfavourable terms of trade in developing countries.
  3. Discuss the importance of balance of payments (BOP) in macroeconomics.
  4. Disadvantages of protection in macroeconomics
  5. Given that a=270,b=0.8,I=150, and G=60.Find the equilibrium level of national income.
1
Expert's answer
2021-05-18T12:32:18-0400

1.

The three methods used in the estimation of national income are:

a) Product method

b) Income Method

c) Expenditure method

 

a) In the product method flow of all the goods and services is measured in order to calculate the national income, It measures the money value of all the final goods and services that are produced in an country during an accounting year. To avoid the problem of double counting only final goods and services are considered while calculating the national income.

b) In Income method flow of factor payment is measured. factor payment is the money value that is paid to the factors of production for their factor services. There are four factors of production namely Land , Labor, Capital, Entrepreneur which are paid factor payment namely Rent, Wages, Interest and profit respectively.

c) In expenditure method flow of expenditure is measured. Expenditure method is the summation of the Consumption expenditure, Government expenditure, Investment expenditure and Net export (export - Import).


2.

reasons for unfavorable terms of trade:

  • high cost ratios - low productivity of production factors
  • backward technology - don't have access to newer technologies
  • primary goods - agrarian economies. still developing from agrarian to industrial
  • high growth in population - high internal demand with low exportable surplus
  • high dependency - lack of import substitution 
  • lack of adaptability - tend to fall back when inflation starts 

 

(3)

Balance of payment examine exports and imports transactions of goods and services for a given period. It therefore aids the government in analyzing the potential of a particular industry export growth and formulate policy to support that growth.


(4)

  • Higher Prices due to tariffs, quotas, exchange rate controls, or regulations used hence they affect the final product price
  • Limited Choice because of restriction of international competition hence there are fewer goods coming into the country
  • Economic Loss.

(5)

Let the consumption function be given by

C = a + bY

Here, a = 270, b = 0.8

Also, I = 150, G = 60


Equilibrium level of national income in a closed economy is given by

Y = C + I + G

Y = a + bY + I + G

Y(1 - 0.8) = 270 + 150 + 60

Y = 480/0.2

  = 2400